# DAMO: Data- and Model-aware Alignment of Multi-modal LLMs

### v1

@@ Abstract
Direct Preference Optimization (DPO) has shown effectiveness in aligning multi-modal large language models (MLLM) with human preferences. However, existing methods exhibit an imbalanced responsiveness to the data of varying hardness, tending to overfit on the easy-to-distinguish data while underfitting on the hard-to-distinguish data. In this paper, we propose Data- and Model-aware DPO (DAMO) to dynamically adjust the optimization process from two key aspects: (1) a data-aware strategy that incorporates data hardness, and (2) a model-aware strategy that integrates real-time model responses. By combining the two strategies, DAMO enables the model to effectively adapt to data with varying levels of hardness. Extensive experiments on five benchmarks demonstrate that DAMO not only significantly enhances the trustworthiness, but also improves the effectiveness over general tasks. For instance, on the Object HalBench, our DAMO-7B reduces response-level and mentioned-level hallucination by 90.0% and 95.3%, respectively, surpassing the performance of GPT-4V. Code is available at: https://github.com/injadlu/DAMO
@@ 1 Introduction
To address this imbalanced responsiveness issue, we propose Data- and Model-aware direct preference optimization (DAMO), which dynamically adapts $\beta$ to both data hardness and model’s responsiveness, enabling adaptively adjust model’s learning behavior based on the inherent preference dynamics and model’s real-time responses. Specifically, we propose two novel mechanisms:
@@
By combining these strategies via element-wise multiplication, DAMO enables real-time adaptation to both data hardness and model responsiveness, demonstrating strong alignment performance across various evaluation benchmarks. Our contributions are summarized as follows:
@@ 3 Approach
In this section, we describe our DAMO in detail. Specifically, we first illustrate our data-aware preference optimization, then we describe our model-aware preference optimization, and finally, we show our combination strategies for robust preference optimization. Our approach algorithm is listed in Algorithm. 1.
@@ 3.3 Combining for Preference Optimization
Algorithm 1 Algorithm of DAMO.
@@ 4 Experiment
In this section, we elaborate on the effectiveness of our Data- and Model-aware Direct Preference Optimization (DAMO). Specifically, we first introduce the details of our experimental settings. Next, we illustrate the ablation studies, and finally, we compare the results with the state-of-the-art methods over various benchmarks.
@@ 4.2 Ablation Studies
In this section, we evaluate the effects of different components of our DAMO. To this end, we utilize the LLaVA-1.5 7B model as the backbone. For clear illustration, we report both the response and the mentioned-level non-hallucination rate on the Object Hallucination benchmark, where the non-hallucination rate is defined as $100\%-$ hallucination rate. The following are detailed illustrations.
Influences of different components. The performance of various components of DAMO is reported in Table 1, where “DPO” refers to Direct Preference Optimization (Rafailov et al., 2024), “MDPO” represents our Model-aware Preference Optimization, “D2PO” denotes Data-aware Preference Optimization, and “DAMO” corresponds to our combined strategy. The experimental results demonstrate: (1) All strategies significantly outperform the baseline method (DPO), especially our final “DAMO”, achieving more than $10\%$ response level performance gains, highlighting the effectiveness of our method; (2) Compared with “MDPO”, the performance gain of “D2PO” is relatively modest, suggesting that the quality of the preference data is already high, which further validates the efficacy of the preference data construction strategy (Yu et al., 2024c).
Table 1: Experimental results of different components of DAMO.
@@
Table 4: Performance comparisons with state-of-the-art methods on different benchmarks. We report non-hallucination rates in different levels including response level (Non-Rsp.) and mentioned-level (Non-Men.) for Object HalBench (Rohrbach et al., 2018). Hall. refers to the Hallucination Rate for MMHal Bench (Sun et al., 2024). The best results of all methods are indicated in bold, and the second best results are underlined. The compared results are sourced from (Yu et al., 2024c; Gu et al., 2024), and the reported results of LLaVA-1.5, DPO, and DAMO are evaluated using GPT-4-turbo-2024-04-09.
@@ 4.3 Comparison with state-of-the-art methods
In this subsection, we compare our method with state-of-the-art methods under three trustworthy benchmarks: Object HalBench, AMBER, and MMHal-Bench, and two general benchmarks: LLaVA-Bench and MM-Vet. We compare our method against baselines from various aspects, including hallucination-specific baselines, preference optimization-based baselines, and the commercial baseline GPT-4V. Our DAMO is adapted to both the LLaVA-1.5 7B and 13B models. The experimental results, along with results of vanilla DPO and the improvements achieved by DAMO over DPO, are listed in Table 4.
From the experimental results, we can observe that: (1) Compared to DPO, our DAMO achieves substantial performance improvements over all compared benchmarks. It’s noteworthy that on MMHal-Bench, it achieves more than 10% improvements for the 13B model, and more than 28% gain for 7B models, which is significant; (2) DAMO achieves new state-of-the-art over various benchmarks. Our DAMO-7B reduces the response-level and mentioned-level hallucination on the generative Object HalBench by 90.9% and 95.3%. Moreover, DAMO-13B achieves 84.3% Accuracy and 88.1 F1 score on the discriminative AMBER benchmark. These attained results surpass those of GPT-4V, fully demonstrating the effectiveness of DAMO.
@@ 6 Conclusion
In this paper, to address the imbalanced responsiveness to data with varying hardness, we propose Data- and Model-aware Direct Preference Optimization (DAMO) to adaptively adjust the model’s learning behavior from two aspects: (1) a data-aware strategy is proposed to dynamically adapt by incorporating the data hardness; and (2) a model-aware strategy is introduced to adaptively adjust by integrating the model’s real-time responsiveness. Extensive experiments on five benchmarks from distinct aspects demonstrate the effectiveness of our method.
@@ A.1 Qualitative analysis
In this section, we provide qualitative analysis between our Data- and Model-aware preference optimization (DAMO) and the DPO method (Rafailov et al., 2024). The case studies are shown in Figure 6 and Figure 7, and we also include the evaluations of GPT-4. From the case studies, we can observe two key points: (1) while both methods generate correct outputs, DPO tends to be more general, whereas DAMO offers richer, more detailed descriptions. This can be attributed to the integration of data hardness, quantified by the CLIP score between the Image and responses, into the optimization process, enabling DAMO to capture and describe more nuanced visual details; and (2) as evidenced by the GPT-4 evaluation, DAMO generates more vivid and immersive visual descriptions, this can be attributed to the enhancement of the responsiveness to the visual content, resulting in a more expressive and contextually aligned representation of the scene.
Figure 6: Qualitative Comparisons between DAMO and vanilla DPO. Data source: Object HalBench.
Figure 7: Qualitative Comparisons between DAMO and vanilla DPO. Data source: Object HalBench.

### v2

# DAMA: Data- and Model-aware Alignment of Multi-modal LLMs

@@ Abstract
Direct Preference Optimization (DPO) has shown effectiveness in aligning multi-modal large language models (MLLM) with human preferences. However, existing methods exhibit an imbalanced responsiveness to the data of varying hardness, tending to overfit on the easy-to-distinguish data while underfitting on the hard-to-distinguish data. In this paper, we propose Data- and Model-aware DPO (DAMA) to dynamically adjust the optimization process from two key aspects: (1) a data-aware strategy that incorporates data hardness, and (2) a model-aware strategy that integrates real-time model responses. By combining the two strategies, DAMA enables the model to effectively adapt to data with varying levels of hardness. Extensive experiments on five benchmarks demonstrate that DAMA not only significantly enhances the trustworthiness, but also improves the effectiveness over general tasks. For instance, on the Object HalBench, our DAMA-7B reduces response-level and mentioned-level hallucination by 90.0% and 95.3%, respectively, surpassing the performance of GPT-4V. Code is available at: https://github.com/injadlu/DAMA
@@ 1 Introduction
To address this imbalanced responsiveness issue, we propose Data- and Model-aware direct preference optimization (DAMA), which dynamically adapts $\beta$ to both data hardness and model’s responsiveness, enabling adaptively adjust model’s learning behavior based on the inherent preference dynamics and model’s real-time responses. Specifically, we propose two novel mechanisms:
@@
By combining these strategies via element-wise multiplication, DAMA enables real-time adaptation to both data hardness and model responsiveness, demonstrating strong alignment performance across various evaluation benchmarks. Our contributions are summarized as follows:
@@ 3 Approach
In this section, we describe our DAMA in detail. Specifically, we first illustrate our data-aware preference optimization, then we describe our model-aware preference optimization, and finally, we show our combination strategies for robust preference optimization. Our approach algorithm is listed in Algorithm. 1.
@@ 3.3 Combining for Preference Optimization
Algorithm 1 Algorithm of DAMA.
@@ 4 Experiment
In this section, we elaborate on the effectiveness of our Data- and Model-aware Direct Preference Optimization (DAMA). Specifically, we first introduce the details of our experimental settings. Next, we illustrate the ablation studies, and finally, we compare the results with the state-of-the-art methods over various benchmarks.
@@ 4.2 Ablation Studies
In this section, we evaluate the effects of different components of our DAMA. To this end, we utilize the LLaVA-1.5 7B model as the backbone. For clear illustration, we report both the response and the mentioned-level non-hallucination rate on the Object Hallucination benchmark, where the non-hallucination rate is defined as $100\%-$ hallucination rate. The following are detailed illustrations.
Influences of different components. The performance of various components of DAMA is reported in Table 1, where “DPO” refers to Direct Preference Optimization (Rafailov et al., 2024), “MDPO” represents our Model-aware Preference Optimization, “D2PO” denotes Data-aware Preference Optimization, and “DAMA” corresponds to our combined strategy. The experimental results demonstrate: (1) All strategies significantly outperform the baseline method (DPO), especially our final “DAMA”, achieving more than $10\%$ response level performance gains, highlighting the effectiveness of our method; (2) Compared with “MDPO”, the performance gain of “D2PO” is relatively modest, suggesting that the quality of the preference data is already high, which further validates the efficacy of the preference data construction strategy (Yu et al., 2024c).
Table 1: Experimental results of different components of DAMA.
@@
Table 4: Performance comparisons with state-of-the-art methods on different benchmarks. We report non-hallucination rates in different levels including response level (Non-Rsp.) and mentioned-level (Non-Men.) for Object HalBench (Rohrbach et al., 2018). Hall. refers to the Hallucination Rate for MMHal Bench (Sun et al., 2024). The best results of all methods are indicated in bold, and the second best results are underlined. The compared results are sourced from (Yu et al., 2024c; Gu et al., 2024), and the reported results of LLaVA-1.5, DPO, and DAMA are evaluated using GPT-4-turbo-2024-04-09.
@@ 4.3 Comparison with state-of-the-art methods
In this subsection, we compare our method with state-of-the-art methods under three trustworthy benchmarks: Object HalBench, AMBER, and MMHal-Bench, and two general benchmarks: LLaVA-Bench and MM-Vet. We compare our method against baselines from various aspects, including hallucination-specific baselines, preference optimization-based baselines, and the commercial baseline GPT-4V. Our DAMA is adapted to both the LLaVA-1.5 7B and 13B models. The experimental results, along with results of vanilla DPO and the improvements achieved by DAMA over DPO, are listed in Table 4.
From the experimental results, we can observe that: (1) Compared to DPO, our DAMA achieves substantial performance improvements over all compared benchmarks. It’s noteworthy that on MMHal-Bench, it achieves more than 10% improvements for the 13B model, and more than 28% gain for 7B models, which is significant; (2) DAMA achieves new state-of-the-art over various benchmarks. Our DAMA-7B reduces the response-level and mentioned-level hallucination on the generative Object HalBench by 90.9% and 95.3%. Moreover, DAMA-13B achieves 84.3% Accuracy and 88.1 F1 score on the discriminative AMBER benchmark. These attained results surpass those of GPT-4V, fully demonstrating the effectiveness of DAMA.
@@ 6 Conclusion
In this paper, to address the imbalanced responsiveness to data with varying hardness, we propose Data- and Model-aware Direct Preference Optimization (DAMA) to adaptively adjust the model’s learning behavior from two aspects: (1) a data-aware strategy is proposed to dynamically adapt by incorporating the data hardness; and (2) a model-aware strategy is introduced to adaptively adjust by integrating the model’s real-time responsiveness. Extensive experiments on five benchmarks from distinct aspects demonstrate the effectiveness of our method.
@@ A.1 Qualitative analysis
In this section, we provide qualitative analysis between our Data- and Model-aware preference optimization (DAMA) and the DPO method (Rafailov et al., 2024). The case studies are shown in Figure 6 and Figure 7, and we also include the evaluations of GPT-4. From the case studies, we can observe two key points: (1) while both methods generate correct outputs, DPO tends to be more general, whereas DAMA offers richer, more detailed descriptions. This can be attributed to the integration of data hardness, quantified by the CLIP score between the Image and responses, into the optimization process, enabling DAMA to capture and describe more nuanced visual details; and (2) as evidenced by the GPT-4 evaluation, DAMA generates more vivid and immersive visual descriptions, this can be attributed to the enhancement of the responsiveness to the visual content, resulting in a more expressive and contextually aligned representation of the scene.
Figure 6: Qualitative Comparisons between DAMA and vanilla DPO. Data source: Object HalBench.
Figure 7: Qualitative Comparisons between DAMA and vanilla DPO. Data source: Object HalBench.